# Effective Interplay between Sparsity and Quantization: From Theory to Practice

### v1

@@ Abstract
The increasing size of deep neural networks necessitates effective model compression to improve computational efficiency and reduce their memory footprint. Sparsity and quantization are two prominent compression methods that have individually demonstrated significant reduction in computational and memory footprints while preserving model accuracy. While effective, the interplay between these two methods remains an open question. In this paper, we investigate the interaction between these two methods and assess whether their combination impacts final model accuracy. We mathematically prove that applying sparsity before quantization is the optimal sequence for these operations, minimizing error in computation. Our empirical studies across a wide range of models, including OPT and Llama model families (125M-8B) and ViT corroborate these theoretical findings. In addition, through rigorous analysis, we demonstrate that sparsity and quantization are not orthogonal; their interaction can significantly harm model accuracy, with quantization error playing a dominant role in this degradation. Our findings extend to the efficient deployment of large models in resource-limited compute platforms and reduce serving cost, offering insights into best practices for applying these compression methods to maximize efficacy without compromising accuracy. 111Code and data are available at: https://bit.ly/quantization-sparsity-interplay
@@ 1 Introduction
Recent breakthroughs in deep neural networks (DNNs) have surpassed human-level capabilities across various tasks such as text generation, machine translation, and computer vision. However, this success comes with significant challenges due to the exponential growth in the size and complexity of DNN models and datasets [3, 69, 52, 58, 1, 57, 25, 46, 37], which complicates their practical deployment and efficient serving. Delivering efficient and real-time inference for these large models is constrained by arithmetic density (throughput/silicon area [11, 6, 20]), memory footprint, and the pressure on memory bandwidth across various hardware platforms (e.g. GPU [45], TPU [43]).
@@
Among various efficiency efforts, model compression has emerged as a crucial solution to effectively address the challenges associated with large models [38, 11, 60, 6, 4, 70, 66, 20, 49, 50, 21, 31, 16, 27, 30, 29] with quantization standing out as a prominent method in terms of overall compression ratio. Quantization effectively reduces the precision of model tensors from native floating-point representation to formats such as FP16 [38], BFloat16 [60], INT8 [59, 68], and others [11, 6, 4, 70, 66, 20, 49, 50].
@@
To further enhance compression, researchers often use sparsity, pruning elements of tensors that are least significant to model accuracy. This reduction in parameters decreases the memory footprint of models and potentially improves computational efficiency by eliminating unnecessary computations [56]. Sparsity methods typically prune elements of a tensor based on their magnitudes (a proxy for the importance of an element to model accuracy) [19, 42, 34, 9, 2]. Magnitude-based sparsity, especially when combined with fine-tuning, offers straightforward implementation and the potential to achieve noticeable compression rates with negligible impact on accuracy [28, 51].
While prior work has extensively studied the impact of sparsity and quantization in isolation, their combined impact is less explored. A few studies [18, 61, 33] suggest that the combination of these two methods does not significantly hurt accuracy. Unfortunately, these studies either primarily focus on CNNs [18] or do not fully investigate the potential combined impact because they exclusively quantize weight tensors while keeping activations in high precision [61, 33].
@@
In this paper, we aim to systematically study the interplay between quantization and sparsity. Quantization and sparsity leverage two fundamentally separate computational properties of DNNs, presenting an opportunity for their combined use. However, their combined impact on model accuracy involves complex dynamics due to the introduction of errors in tensors. We hypothesize that the interplay between quantization and sparsity can result in additional errors, based on the following insights. First, applying quantization before sparsity (QS) may adversely disrupt the relative importance of tensor elements, potentially leading to the removal of significant elements of a tensor. Second, applying sparsity before quantization (SQ) can introduce additional errors in dot product calculations, as these are influenced by the magnitudes and precision of the involved elements. To the best of our knowledge, we are the first to study the interplay between quantization and sparsity in depth, to identify the conditions under which accuracy can be preserved or compromised. Our contributions are summarized below:
Theoretical Analysis of Sparsity-Quantization Interplay: We provide the first in-depth theoretical analysis of how sparsity and quantization interact. Our mathematical proofs establish that applying sparsity before quantization (S Q) is the optimal sequence for compression.
Non-orthogonality and Error Dynamics of Compression Methods: We mathematically demonstrate that sparsity and quantization are not orthogonal operations. Combining them introduces additional errors beyond the sum of their individual errors. We analyze the specific error dynamics arising from this interaction, highlighting the dominant role of quantization error in the overall degradation of model accuracy.
@@
Extensive Empirical Study: We validate our theoretical findings through extensive experiments covering a diverse range of models, including prominent LLMs (OPT, LLaMA) and ViTs. These experiments support our hypotheses about the optimal order of compression operations and underscore the non-orthogonality of sparsity and quantization, affirming the practical relevance of our theoretical analysis. Our experiments demonstrate that applying quantization and sparsity in a sub-optimal order can cause up to 7.96 point increase in perplexity.
@@ 2 Related work
Quantization. The ever-growing size of DNN models has spurred extensive research into using narrow numerical formats for inference to reduce memory footprint and improve computational efficiency [38, 11, 60, 6, 4, 70, 66, 20, 49, 50]. Numerical formats employ scaling factors to adjust their dynamic range and can be categorized based on the granularity and their levels. Table 3 in AppendixA summarizes the most common numerical formats.
Sparsity. Sparsity methods [21, 31, 16, 27, 30] aim to reduce computational and storage overhead in DNNs by selectively eliminating tensor elements according to various sparsity mask selection criteria. Broadly, these methods fall into two main categories based on sparsity patterns: unstructured [18, 17, 13, 12] and structured [62, 65, 26, 42, 48, 71, 54, 24, 34].
@@
Combining quantization and sparsity. Prior work has studied the combination of quantization and sparsity and its impact on model performance: sparsity followed by quantization [67, 15, 14, 47, 42, 32, 18] and quantization followed by sparsity [23, 22, 63, 42]. There are two missing pieces of information from these pieces of prior work. First, consensus on the optimal order of compression operations is not clear. Some favor one order without clear justification, while others treat the methods as independent compression schemes. Second, there is lack of theoretical grasp on how sparsity and quantization errors interact and influence final model performance. Appendix A covers additional related work.
3 Mathematical analysis of orthogonality of quantization and sparsity
@@
This section provides a mathematical analysis of the interplay between quantization and sparsity, formalizing these compression methods and examining their composition at both the tensor and dot product levels. Our analysis centers on quantization methods that reduce the bit-width of model weights and activations using block-wise numerical formats, which are prevalent in practical implementations [6, 11, 70, 49, 50, 40]. These formats determine the scaling factor based on the maximum element within a block of elements. We denote a quantization method employing these numerical formats as max-scaled block-wise quantization. We use magnitude-based sparsity for both unstructured and N:M structured sparsity.
@@
Let be a block of numbers and denotes the quantization bit-width. Max-scaled block-wise quantization is a transformation of the block such that
where is the scaling factor. quantizes the given element with the scaling factor and the number of mantissa bits . The exact form of depends on the numerical format and can be found in Appendix J. For instance, INT quantization transformation is defined as follows:
@@
where is the -th largest element in . The same formula can be adjusted to represent unstructured sparsity by defining as the -th largest element in the tensor, where , is the number of elements in the tensor, and is the operation of rounding to the nearest integer.
In the remaining of this section, we delve into the composition of quantization and sparsity at two different levels. First, we examine the effects of applying this composition in different orders at the tensor level, observing how individual tensors are altered. Then, we explore how the composition influences the result of the dot product operation.
@@ 3.1 Tensor-level analysis
Quantization and sparsity transformations inherently introduce errors by decreasing precision or pruning tensor elements. To study the composition of quantization and sparsity transformations at the tensor level, we introduce formal definitions of transformation error and orthogonality. We prove that orthogonality between sparsity and quantization does not persist within this composition. The following definition formalizes the error for a specific transformation at the block level, which consists of a subset of tensor elements.
@@
Combining two compression methods (transformations) is expected to introduce additional error. Any error introduced by the first transformation becomes part of the input to the second transformation, potentially amplifying the initial error and resulting in a larger overall error.
@@
We define two transformations and to be orthogonal if any order of their composition does not introduce any additional error, and thus, the following inequalities hold:
@@
Let be the max-scaled block-wise quantization and be the magnitude-based sparsity transformation. Applying sparsity before quantization does not introduce additional error:
@@
Proof.
Let represent the number of elements pruned from the block by the sparsity transformation. Without loss of generality, we assume that the last elements in the block are pruned, as permuting the elements does not affect the block’s norm. Since the sparsity transformation does not prune the largest element in the block, the parameter of quantization remains unchanged. Consequently, the quantization error for the non-zero components before and after sparsity remains the same:
To demonstrate that equality is attainable, we consider the block of floating-point numbers , INT4 quantization and sparsity.
The L1-norms of the transformation errors are the following:
Therefore, is attainable. ∎
@@
Proof.
Thus, for this particular input , the inequality holds true. ∎
Moreover, a global upper bound exists for the additional error arising from this specific order of transformations (quantizationsparsity). This upper bound is solely determined by the quantization method and the parameters of the sparsity scheme, independent of the input data. The following theorem precisely quantifies the magnitude of this additional error.
@@
Proof of Theorem 3.7 can be found in Appendix I. As a corollary of Theorem 3.5, 3.6, and 3.7, it follows that the optimal order of transformations is sparsity followed by quantization, as this sequence does not introduce any additional error. Moreover, according to Definition 3.4, quantization and sparsity are not orthogonal at the tensor level.
3.2 Orthogonality of dot product operation
In this section, we delve into the error linked with the dot product operation, which is the primary operation in DNNs. Our analysis focuses on scenarios where weight tensors undergo sparsification and quantization, while activation tensors solely undergo quantization. We first extend the definition of transformation error to the dot product level.
@@ 3.2 Orthogonality of dot product operation
At the dot product level, we define two compression methods as orthogonal if any order of their composition does not introduce additional error, akin to Definition 3.4. In the following theorem, we demonstrate that any combination of sparsity and quantization yields additional error, rendering these two methods non-orthogonal.
@@
Let be the max-scaled block-wise quantization, be the magnitude-based sparsity transformation, be the composition which is either or and be the identity function. Max-scaled quantization and sparsity produce additional error in any order, given that only the second operand, i.e., weight, is sparsified:
@@
where is the correction error vector of the composition:
Proof.
Error of the composition can be written as the following:
After adding the norms, we obtain the following:
where and are the additional error terms.
To prove non-orthogonality, consider the blocks of floating-point numbers and , HBFP4 quantization and 1:2 sparsity . For these particular values of and , the inequality: holds true. The exact calculations can be found in Appendix I. ∎
Analysis of the additional error. As a corollary of Theorem 3.9, the combination of max-scaled quantization and sparsity is not orthogonal, resulting in two additional error terms.
The term encompasses the correction vector of the composition , which propagates the error from the tensor level to the dot product level. Depending on the order of the composition the value of varies.
However, if quantization is applied first, certain elements in the block may become equal, resulting in the sparsity removing a different set of elements. Formally, if is pruned by but not by , there exists a , where , such that . In this scenario, and . Otherwise, it only contains the quantization errors of the pruned elements. Therefore, the magnitude of the correction vector for the composition is generally larger than that of the reverse order. Apart from the quantization errors, the correction vector also contains pruned elements, which are larger by orders of magnitude except in a few improbable edge cases. This results in the overall value of being larger, implying that the order of applying sparsity first and then quantization is optimal for dot product.
The term also contributes to the additional error, encoding the interaction between the error vectors and . However, is less significant than , as it contains the quantization error, the norm of which is generally orders of magnitude lower than the norm of the original block. In addition, contains the sparsity error, which involves the smallest weights, diminishing the significance of this additional error.
Finally, to experimentally validate our theoretical findings, we define orthogonality bound to assess whether the transformations are orthogonal.
Definition 3.10 (Orthogonality bound).
Let be a DNN model under consideration, be an evaluation metric that measures the performance of the model (e.g. Perplexity or Cross-Entropy Loss), be the evaluation metric of the model with transformation , which is either sparsity or quantization . Moreover, let be the evaluation metric error of the transformation for the model . We define orthogonality bound as:
If the compression methods are not orthogonal, and the evaluation metric indicates better model performance with lower values, we expect the compressed model’s evaluation metric to exceed the orthogonality bound.

### v2

@@ Abstract
The increasing size of deep neural networks (DNNs) necessitates effective model compression to reduce their computational and memory footprints. Sparsity and quantization are two prominent compression methods that have been shown to reduce DNNs’ computational and memory footprints significantly while preserving model accuracy. However, how these two methods interact when combined together remains a key question for developers, as many tacitly assume that they are orthogonal, meaning that their combined use does not introduce additional errors beyond those introduced by each method independently. In this paper, we provide the first mathematical proof that sparsity and quantization are non-orthogonal. We corroborate these results with experiments spanning a range of large language models, including the OPT and LLaMA model families (with 125M to 8B parameters), and vision models like ViT and ResNet. We show that the order in which we apply these methods matters because applying quantization before sparsity may disrupt the relative importance of tensor elements, which may inadvertently remove significant elements from a tensor. More importantly, we show that even if applied in the correct order, the compounded errors from sparsity and quantization can significantly harm accuracy. Our findings extend to the efficient deployment of large models in resource-constrained compute platforms to reduce serving cost, offering insights into best practices for applying these compression methods to maximize hardware resource efficiency without compromising accuracy. 111Code and data are available at: https://bit.ly/quantization-sparsity-interplay
@@ 1 Introduction
Recent breakthroughs in deep neural networks (DNNs) have surpassed human-level capabilities across various tasks such as text generation, machine translation, and computer vision. Unfortunately, this achievement is accompanied by significant challenges due to the exponential growth in the size and complexity of DNN models and datasets (Brown et al., 2020; Zhang et al., 2022b; Scao et al., 2022; Touvron et al., 2023; Almazrouei et al., 2023; Anil et al., 2023; Jiang et al., 2023; OpenAI, 2023; Mesnard et al., 2024), which complicates their practical deployment and efficient serving. Delivering efficient and real-time inference for these large models is constrained by arithmetic density (throughput/silicon area (Drumond et al., 2018a; Darvish Rouhani et al., 2020a; Harma et al., 2022)), memory footprint, and the pressure on memory bandwidth across various hardware platforms (e.g. GPU (Nvidia, 2022), TPU (Google, 2023)).
@@
Among various efficiency efforts, model compression has emerged as a crucial solution to effectively address the challenges associated with large models (Micikevicius et al., 2018; Drumond et al., 2018b; a; Wang & Kanwar, 2019; Darvish Rouhani et al., 2020b; a; Dai et al., 2021; Zhang et al., 2022a; Yeh et al., 2022; Harma et al., 2022; Rouhani et al., 2023a; b; Hassibi et al., 1993; LeCun et al., 1989; Frantar et al., 2023; Kao et al., 2022; Lasby et al., 2023; Kuzmin et al., 2023) with quantization standing out as a prominent method in terms of overall compression ratio achieved. Quantization effectively reduces the precision of model tensors from native floating-point representation to formats such as FP16 (Micikevicius et al., 2018), BFloat16 (Wang & Kanwar, 2019), and INT8 (van Baalen et al., 2023; Zafrir et al., 2019). In recent years, block-wise numerical formats have gained prominence (Drumond et al., 2018b; a; Darvish Rouhani et al., 2020b; a; Dai et al., 2021; Zhang et al., 2022a; Yeh et al., 2022; Harma et al., 2022; Rouhani et al., 2023a; b), particularly due to their ability to reduce memory footprint and increase arithmetic density in sub-8-bit regimes, while preserving accuracy with minimal hardware overhead.
@@
Beyond quantization, researchers are also exploiting sparsity to further compress models, pruning elements of tensors that are least significant to preserve model accuracy. This reduction in the number of parameters decreases the models’ memory footprint and eliminates potentially unnecessary computation (Sze et al., 2020). The most popular sparsity method is magnitude-based sparsity which prunes elements of a tensor based on their magnitudes (i.e., a proxy for the importance of an element to model accuracy) (Han et al., 2015b; Mishra et al., 2021b; Lu et al., 2023; Ding et al., 2023; Bambhaniya et al., 2024), which when combined with fine-tuning achieves noticeable compression rates with negligible impact on accuracy (Kurtic et al., 2022; Sanh et al., 2020).
While combining sparsity and quantization provides significant gains in arithmetic density and memory footprint, it may inadvertently have a high impact on model accuracy. Prior work tacitly assumes that these two methods are orthogonal, meaning that their combined use does not introduce additional errors beyond those of each method individually. These include studies focusing on CNNs (Han et al., 2015a; Wang et al., 2020), which are more resilient to quantization errors due to the absence of dot-product outliers in activation tensors (Xiao et al., 2023). Because quantization error in CNNs is relatively low, additional errors introduced by combining the two are minimal. Other studies focus only on compressing weights without quantizing activations (Li et al., 2020; Liu et al., 2023), which also leads to low overall error when combined with sparsity. These studies fall short of properly investigating the combined impact of these two compression methods to maximize hardware resource efficiency without compromising accuracy.
@@
In this paper, we study the interplay between sparsity and quantization systematically. Sparsity and quantization leverage fundamentally separate computational properties of DNNs, but their combined impact on model accuracy involves complex interactions due to the introduction of errors in tensors. We hypothesize that sparsity and quantization are non-orthogonal based on the following two insights. First, applying quantization before sparsity (QS) may adversely disrupt the relative importance of tensor elements, leading to the removal of significant elements of a tensor with a significant impact on model accuracy. Second, applying sparsity before quantization (SQ) can introduce additional errors in dot product calculations, as these are influenced by the magnitudes and precision of the involved elements requiring a careful investigation. To the best of our knowledge, we are the first to study the interplay between sparsity and quantization in depth to identify the conditions under which accuracy can be preserved or compromised. Our contributions are summarized below:
Non-Orthogonality of Sparsity and Quantization: We prove mathematically that sparsity and quantization are non-orthogonal operations. Our per-layer error analysis shows that combination of the two introduces compounded errors and a degradation of model accuracy. Our findings challenge the conventional wisdom that these methods can be combined without a significant impact on accuracy.
Corroborating the Optimal Compression Order: Although applying sparsity before quantization (SQ) is the most commonly adopted approach, the optimal order has not been formally demonstrated in the literature. We provide the first mathematical proof that applying sparsity before quantization (SQ) is optimal. Moreover, we derive the upper bound for the error caused by the sub-optimal order of the transformations at the tensor level. We show that it depends linearly on the number of elements in the tensor and the size of the quantization bin.
@@
Validating Non-orthogonality Empirically: We validate our mathematical findings with experiments covering a diverse range of models, including prominent LLMs (OPT, LLaMA), ViT, and ResNet. These experiments support our hypotheses and mathematical findings, underscoring the non-orthogonality of sparsity and quantization, and the optimal order of compression methods. Our experiments demonstrate that combining sparsity and quantization, even in the optimal order, can cause up to 13% additional error in perplexity.
@@ 2 Related work
Quantization. The ever-growing size of DNN models has spurred extensive research into using narrow numerical formats for inference to reduce memory footprint and improve computational efficiency (Micikevicius et al., 2018; Drumond et al., 2018b; a; Wang & Kanwar, 2019; Darvish Rouhani et al., 2020b; a; Dai et al., 2021; Zhang et al., 2022a; Yeh et al., 2022; Harma et al., 2022; Rouhani et al., 2023a; b). Numerical formats employ scaling factors to adjust their dynamic range and can be categorized based on the granularity and levels of the scaling factors.
Element-wise scaling formats, such as FP32, BFloat16 (Wang & Kanwar, 2019), and FP16 (Micikevicius et al., 2018), consist of sign, mantissa, and exponent components, differing in the bit allocation for each component. Conversely, block-wise scaling formats assign scaling factors to blocks of elements, with block sizes varying by format. For instance, INT8 employs per-tensor scaling, where a single scaling factor is shared by around 1K elements. Recent research highlights the effectiveness of max-scaled fine-grained block-wise scaling formats with block sizes smaller than 100 elements, especially in the sub-8-bit regime for both training and inference (Drumond et al., 2018a; Rouhani et al., 2023a; b; Zhang et al., 2022a; Darvish Rouhani et al., 2020a). Additionally, modern hardware platforms have adopted these techniques. For instance, the upcoming NVIDIA Blackwell GPUs (Nvidia, 2024) will support MXFP.
Sparsity. Sparsity methods (Hassibi et al., 1993; LeCun et al., 1989; Frantar et al., 2023; Kao et al., 2022; Lasby et al., 2023) aim to reduce computational and memory footprints in DNNs by selectively pruning tensor elements according to various sparsity mask selection criteria. Broadly, these methods fall into two main categories based on sparsity patterns: unstructured (Han et al., 2015a; Guo et al., 2016; Frankle & Carbin, 2019; Evci et al., 2020a) and structured (Wen et al., 2016; Yao et al., 2019; Kang, 2020; Mishra et al., 2021b; Pool & Yu, 2021; Zhou et al., 2021; Sun et al., 2021; Hubara et al., 2021; Lu et al., 2023). Unstructured sparsity (Han et al., 2015a; Guo et al., 2016; Frankle & Carbin, 2019; Evci et al., 2020a) involves removing individual tensor elements without any specific pattern. Structured sparsity (Wen et al., 2016), on the other hand, employs specific patterns when pruning tensor elements. Recent work (Yao et al., 2019; Kang, 2020) has highlighted the effectiveness of fine-grained N:M structured sparsity in mitigating model accuracy loss.
The fundamental operation in any sparsification scheme is selecting candidate elements for pruning among which magnitude-based sparsity (Han et al., 2015b) is one of the most widely used methods (Lu et al., 2023; Ding et al., 2023; Bambhaniya et al., 2024). In addition, recent work has introduced one-shot pruning methods, such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023), aiming to eliminate the need for an additional fine-tuning phase. However, evidence suggests that incorporating fine-tuning still improves accuracies significantly (Sun et al., 2023; Syed et al., 2023; Lu et al., 2024).
@@
Combining sparsity and quantization. Prior work has studied the combination of sparsity and quantization and its impact on model accuracy in both orders: sparsity followed by quantization (Yu et al., 2023; Frantar & Alistarh, 2023; 2022; Park et al., 2022; Mishra et al., 2021a; Li et al., 2020; Han et al., 2015a) and quantization followed by sparsity (Hu et al., 2021; Hawks et al., 2021; Wu et al., 2023; Mishra et al., 2021b). There are two missing pieces of information from prior work. First, consensus on the optimal order of compression operations is not established. A few studies raise the question and experiment with both orders to determine the best approach (Wu et al., 2023; Wang et al., 2022; Zandonati et al., 2023; Park et al., 2019; Yu et al., 2020; Mishra et al., 2021a; Kozlov et al., 2021; Zhang et al., 2021), while others treat the methods as orthogonal compression schemes. Second, there is a lack of mathematical grasp on how sparsity and quantization errors interact and influence final model performance.
3 Non-orthogonality of sparsity and quantization
@@
This section provides a mathematical analysis of the interplay between sparsity and quantization, formalizing these compression methods and examining their combination, henceforth referred to as (mathematical) composition, at both the tensor and dot-product levels. For the remainder of the paper, we use the following notions: (1) Tensor level refers to structures that encompass both weight and activation tensors; and (2) Dot-product level pertains to the computation of inner products within these tensors, such as the matrix multiplication operation between weights and activations during the forward pass. Our analysis centers on quantization methods that reduce the bit-width of model weights and activations using block-wise numerical formats, which are prevalent in practical implementations (Darvish Rouhani et al., 2020a; Drumond et al., 2018a; Zhang et al., 2022a; Rouhani et al., 2023a; b; Micikevicius et al., 2022). These formats determine the scaling factor based on the element with the maximum magnitude within the block. We refer to any quantization method employing these numerical formats as max-scaled block-wise quantization. We use magnitude-based sparsity for both unstructured and N:M structured sparsity.
@@
Let be a block of numbers and denote the quantization bit-width. Max-scaled block-wise quantization is a transformation of the block such that
where is the scaling factor. quantizes the given element with the scaling factor and the number of mantissa bits . The exact form of depends on the numerical format and can be found in Appendix K. For instance, INT quantization transformation is defined as follows:
@@
where is the -th largest element in the set . The same formula can be adjusted to represent unstructured sparsity by defining as the -th largest element in the tensor, where , is the number of elements in the tensor, and is the operation of rounding to the nearest integer.
In the remainder of this section, we delve into the composition of sparsity and quantization at two different levels. First, we examine the effects of applying this composition in different orders at the tensor level, observing how individual tensors are altered. Then, we explore how the composition influences the result of the dot product operation.
@@ 3.1 Tensor-level analysis
Sparsity and quantization transformations inherently introduce errors by decreasing precision or pruning tensor elements. To study the composition of sparsity and quantization transformations at the tensor level, we introduce formal definitions of transformation error and orthogonality in compression. We prove that orthogonality in compression between sparsity and quantization does not persist within this composition.
The following definition formalizes the error for a specific transformation at the block level, which consists of a subset of tensor elements.
@@
Composing two compression methods (transformations) is expected to introduce additional errors. Any error introduced by the first transformation becomes part of the input to the second transformation, potentially amplifying the initial error and resulting in a larger overall error.
@@
We define two transformations and to be orthogonal in compression222In the remainder of the paper, we use the term ”orthogonal” to refer to ”orthogonal in compression” for simplicity. if any order of their composition does not introduce any additional error, and thus, the following inequalities hold:
where is an norm, .
@@
Let be the max-scaled block-wise quantization and be the magnitude-based sparsity transformation. Applying sparsity before quantization does not introduce any additional error:
@@
The proof of Theorem 3.5 can be found in Appendix J. The main idea behind the proof is that as the sparsity transformation does not prune the largest element in the block, the quantization parameter remains unchanged. Consequently, the quantization error for the non-zero components before and after sparsity remains the same.
@@
Moreover, a global upper bound exists for the additional error arising from this specific order of transformations (QS). This upper bound is solely determined by the quantization method and the parameters of the sparsity type, independent of the input data. The following theorem precisely quantifies the magnitude of this additional error.
@@
The general formulation of Theorem 3.7 for all norms and the proof of Theorem 3.6 and 3.7 can be found in Appendix J. As a corollary of Theorem 3.5, 3.6, and 3.7, it follows that the optimal order of transformations is sparsity followed by quantization, as this sequence does not introduce any additional error. Moreover, according to Definition 3.4, sparsity and quantization are non-orthogonal at the tensor level.
3.2 Dot-product-level analysis
In this section, we delve into the error linked with the dot product operation, which is the primary operation in DNNs. Our analysis focuses on scenarios where weight tensors undergo sparsity and quantization, while activation tensors solely undergo quantization. We first extend the definition of transformation error to the dot-product level.
@@ 3.2 Dot-product-level analysis
At the dot-product level, we define two compression methods as orthogonal if their composition, in any order, does not introduce additional error, akin to Definition 3.4.
Definition 3.9 (Dot-product-level orthogonality).
Let denote the inputs of transformations and , and the dot product operation . Let the transformation be applied to both and , and transformation be applied only to . Let denote a composition of and in any order, . We define two transformations and to be orthogonal on the dot-product level if any order of their composition applied to the second term does not introduce any additional error:
In the following theorem, we demonstrate that any composition of sparsity and quantization yields additional error, rendering these two methods non-orthogonal.
@@
Let be the max-scaled block-wise quantization, be the magnitude-based sparsity transformation, be the composition which is either or and be the identity function. Composition of max-scaled quantization and sparsity in any order produces additional error in any order, given that only the second operand, i.e., weight, is pruned:
@@
where is defined as the correction error vector of the composition:
Analysis of the additional error. As a corollary of Theorem 3.10, the composition of max-scaled sparsity and quantization is non-orthogonal, resulting in two additional error terms.
The term incorporates the correction vector of the composition , which carries the additional error from the tensor level to the dot-product level. Depending on the order of the composition, the value of varies. When quantization precedes sparsity, certain elements within a block may become equal due to quantization, leading the sparsity step to prune different elements than it would on the original tensor. This introduces additional error, as previously significant elements may be inadvertently pruned. If sparsity precedes quantization, the correction vector exclusively comprises the quantization errors of the pruned elements. As a result, the magnitude of the additional error is typically smaller compared to the reverse order.
The term also contributes to the additional error, encoding the interaction between the error vectors and . However, since the norm of quantization and sparsity errors are generally smaller than the norm of the quantization block, this term is less significant than .
We provide a more detailed explanation of the additional error analysis in Appendix M
Finally, to experimentally validate our mathematical findings, we define a metric, orthogonality threshold to assess whether the transformations are orthogonal.
Definition 3.11 (Orthogonality threshold).
Let be a DNN model under consideration, be an evaluation metric that measures the performance of the model (e.g., perplexity or cross-entropy loss), be the evaluation metric of the model with transformation , which is either sparsity or quantization . Moreover, let be the evaluation metric error of the transformation for the model . We define orthogonality threshold as:
If the compression methods are non-orthogonal, and the evaluation metric improves with lower values (e.g., perplexity), we expect the compressed model’s evaluation metric (e.g., perplexity) to worsen and thus exceed the orthogonality threshold due to compounded errors. Similarly, if the compression methods are non-orthogonal, and the evaluation metric improves with higher values (e.g., accuracy), we expect the compressed model’s evaluation metric (e.g., accuracy) to decrease and thus fall below the orthogonality threshold.